Use the counting techniques from the last chapter. A bag contains two red marbles, four green ones, one fluorescent pink one, two yellow ones, and two orange ones. Suzan grabs three at random. Find the probability of the indicated event.
She gets all the red ones, given that she gets the fluorescent pink one.
I don't understand how to set up this problem!

Answer by stanbon(75887)

Use the counting techniques from the last chapter. A bag contains two red marbles, four green ones, one fluorescent pink one, two yellow ones, and two orange ones. Suzan grabs three at random. Find the probability of the indicated event.
She gets all the red ones, given that she gets the fluorescent pink one.
-------
P(2red | pink) = P(2red and pink)/P(pink)
---------------------------
# of ways to get 2red and pink = 1
# of possible triples: 12C3 = 220
---
P(2red and pink) = 1/220
---
P(pink) =
# of ways to get 1 pink = 1*11*10
# of possible triples = 220
P(1 pink in 3) = 110/220 = 1/2
-------
P(2red |pink) = (1/220)/(1/2) = 2/220 = 1/110
----------------------------------------------
Cheers,
Stan H.
================